The following table of data were obtained from a baseball statistics source, showing hit information for two well known baseball players. Suppose that one hit from the table is randomly selected. Single Double Triple Home Run Totals Henry Ruth 1834 624 161 495 3114 Babe Aaron 1996 661 127 479 3263 Totals 3830 1285 288 974 6377 a. (396) Complete the above table. b. (396) P( Single ) С. (3%) P( Babe Aaron ) d. (3%) P( Henry Ruth and Triple e. (3%) P( Babe Aaron I Double)-
media%2Fbd9%2Fbd9cf675-f35b-470e-9786-26
media%2F518%2F518ad0ae-d8ed-405a-886e-11
media%2Fa75%2Fa750350d-1f19-4a60-a6d3-16
media%2Fbf9%2Fbf9690fe-83da-4bf9-a644-b8
NE8823 2
dt
Answer #1

Ans:

a)

Single Double triple Home Run Totals
Henry 1834 624 161 495 3114
Babe 1996 661 127 479 3263
Totals 3830 1285 288 974 6377

b)

P(Single)=3830/6377=0.601

c)

P(Babe Aron)=3263/6377=0.512

d)

P(Henry ruth and triple)=161/6377=0.025

e)

P(Babe Aron/Double)=661/1285=0.514